Q. Choose the correct answer
Q. Which of the following differential equation (with dependent variable y) are linear
(i) (x+y)dy=x2dx+y2dy
(ii) (x2+y)dy=x2dx+ydy
(iii)(x2+y)dy=sin ydx
1
Expert's answer
2019-07-08T10:27:45-0400

(II)


First order linear differential equation has the form


"\frac{dy}{dx}+P(x)y=Q(x)"

Dividing the giving equations by dx, we can have the beforementioned form only in the 2nd case


"(x^2+y)\frac{dy}{dx}=x^2+y\frac{dy}{dx}"

"x^2\frac{dy}{dx}+y\frac{dy}{dx}-y\frac{dy}{dx}=x^2""x^2\frac{dy}{dx}=x^2"

"\frac{dy}{dx}=1"